# Attention-Guided Perturbation for Unsupervised Image Anomaly Detection

### v2

@@ 4.2 Anomaly Detection on MVTec-AD
Bottle 99.2 / 99.1 / - 100.0 / 98.7 / 96.6 100.0 / 98.6 / - - / 99.2 / - 100.0 / 98.0 / - 100.0 / 99.1 / 94.6
Cable 99.9 / 97.6 / - 95.0 / 97.4 / 91.0 99.5 / 98.4 / - - / 97.3 / - 94.8 / 96.8 / - 100.0 / 98.7 / 93.9
Capsule 98.5 / 94.3 / - 96.3 / 98.7 / 95.8 98.1 / 98.8 / - - / 99.1 / - 97.7 / 98.9 / - 97.9 / 98.3 / 90.4
Hazelnut 100.0 / 99.7 / - 99.9 / 98.9 / 95.5 100.0 / 98.7 / - - / 99.6 / - 100.0 / 97.9 / - 100.0 / 99.0 / 94.4
Metal_nut 98.7 / 99.5 / - 100.0 / 97.3 / 92.3 100.0 / 98.4 / - - / 98.6 / - 100.0 / 98.8 / - 100.0 / 97.0 / 91.5
Pill 98.9 / 97.6 / - 96.6 / 98.2 / 96.4 96.6 / 97.4 / - - / 98.7 / - 99.0 / 98.6 / - 98.2 / 98.6 / 96.4
Screw 93.9 / 97.6 / - 97.0 / 99.6 / 98.2 98.1 / 99.4 / - - / 98.5 / - 98.2 / 99.3 / - 97.2 / 99.3 / 96.7
Toothbrush 100.0 / 98.1 / - 99.5 / 99.1 / 94.5 100.0 / 98.7 / - - / 99.3 / - 99.7 / 98.5 97.2 / 98.9 / 90.3
Transistor 93.1 / 90.9 / - 96.7 / 92.5 / 78.0 100.0 / 96.3 / - - / 89.1 / - 100.0 / 97.6 / - 99.5 / 97.3 / 91.3
Zipper 100.0 / 98.8 / - 98.5 / 98.2 / 95.4 99.4 / 98.8 / - - / 99.1 / - 99.9 / 98.9 / - 99.3 / 97.4 / 93.0
Carpet 97.0 / 95.5 / - 98.9 / 98.9 / 97.0 98.7 / 99.0 / - - / 96.1 / - 99.7 / 98.2 / - 100.0 / 99.2 / 95.5
Grid 99.9 / 99.7 / - 100.0 / 99.3 / 97.6 98.2 / 98.7 / - - / 99.1 / - 99.7 / 98.8 / - 100.0 / 98.9 / 96.2
Leather 100.0 / 98.6 / - 100.0 / 99.4 / 99.1 100.0 / 99.3 / - - / 99.7 / - 100.0 / 99.2 / - 100.0 / 99.5 / 97.7
Tile 99.6 / 99.2 / - 99.3 / 95.6 / 90.6 98.7 / 95.6 / - - / 98.0 / - 99.8 / 97.0 / - 100.0 / 97.5 / 89.1
Wood 99.1 / 96.4 / - 99.2 / 95.3 / 90.9 99.2 / 95.0 / - - / 97.7 / - 100.0 / 94.5 / - 98.8 / 96.7 / 88.5
Mean 98.0 / 97.3 / - 98.5 / 97.8 / 93.9 99.1 / 98.1 / 93.5 98.6 / 97.9 / - 99.6 / 98.1 / - 99.2 / 98.3 / 93.1
Evaluation on MVTec-AD under the one-class setting. As shown in Table 2, we can observe that compared with the two synthetic data-based methods, DRAEM Zavrtanik et al. (2021) and DeSTSeg Roth et al. (2022), we outperform them by a large margin. Specifically, compared with DRAEM, we outperform it by 1.2% and 1.0% in I-AUC and P-AUC, respectively. And compared to DeSTSeg, we also bring gains of 0.6% I-AUC and 0.4% P-AUC. Moreover, we achieved the best localization result of 98.3% P-AUC, while the detection performance is slightly lower than SimpleNet, ranking second among previous arts. It should be noted that when SimpleNet is adapted to the multi-class setup, its performance significantly drops, as evidenced in Table 1. In contrast, our method demonstrates robust performance across both setups, highlighting its strong generalization capability and suitability for diverse real-world applications.
Table 3: Anomaly detection results on VisA. denotes the reproduced results using the official code. Bold and underline indicate the best and the second best, respectively.
Category
DRÆM*
(ICCV 2021)
UniAD*
(NeurIPS 2022)
SimpleNet*
(CVPR 2023)
DiAD
(AAAI 2024)
ViTAD
(arXiv 2023)
Ours
I-AUC / P-AUC / PRO (%)
PCB1 71.9 / 94.6 / 52.9 92.8 / 93.3 / 64.1 91.6 / 99.2 / 83.6 88.1 / 98.7 / 80.2 95.8 / 99.5 / 89.6 97.1 / 99.5 / 92.2
PCB2 78.4 / 92.3 / 66.2 87.8 / 93.9 / 66.9 92.4 / 96.6 / 85.7 91.4 / 95.2 / 67.0 90.6 / 97.9 / 82.0 94.7 / 98.0 / 82.0
PCB3 76.6 / 90.8 / 43.0 78.6 / 97.3 / 70.6 89.1 / 97.2 / 85.1 86.2 / 96.7 / 68.9 90.9 / 98.2 / 88.0 92.1 / 98.3 / 81.4
PCB4 97.3 / 94.4 / 75.7 98.8 / 94.9 / 72.3 97.0 / 93.9 / 61.1 99.6 / 97.0 / 85.0 99.1 / 99.1 / 91.8 99.2 / 98.7 / 90.0
Macaroni1 69.8 / 95.0 / 67.0 79.9 / 97.4 / 84.0 85.9 / 98.9 / 92.0 85.7 / 94.1 / 68.5 85.8 / 98.5 / 89.2 91.0 / 98.1 / 86.4
Macaroni2 59.4 / 94.6 / 65.3 71.6 / 95.2 / 76.6 68.3 / 93.2 / 77.8 62.5 / 93.6 / 73.1 79.1 / 98.1 / 87.2 79.8 / 96.9 / 85.0
Capsules 83.4 / 97.1 / 62.9 55.6 / 88.7 / 43.7 74.1 / 97.1 / 73.7 58.2 / 97.3 / 77.9 79.2 / 98.2 / 75.1 78.2 / 98.5 / 78.5
Candles 69.3 / 82.2 / 65.6 94.1 / 98.5 / 91.6 84.1 / 97.6 / 87.6 92.8 / 97.3 / 89.4 90.4 / 96.2 / 85.2 93.6 / 97.7 / 88.5
Cashew 81.7 / 80.7 / 38.5 92.8 / 98.6 / 87.9 88.0 / 98.9 / 84.1 91.5 / 90.9 / 61.8 87.8 / 98.5 / 78.8 94.1 / 99.2 / 87.7
Chewing gum 93.7 / 91.0 / 41.0 96.3 / 98.8 / 81.3 96.4 / 97.9 / 78.3 99.1 / 94.7 / 59.5 94.9 / 97.8 / 71.5 96.1 / 98.9 / 73.5
Fryum 89.1 / 92.4 / 69.5 83.0 / 95.9 / 76.2 88.4 / 93.0 / 85.1 89.8 / 97.6 / 81.3 94.3 / 97.5 / 87.8 93.7 / 97.5 / 86.8
Pipe fryum 82.8 / 91.1 / 61.9 94.7 / 98.9 / 91.5 90.8 / 98.5 / 83.0 96.2 / 99.4 / 89.9 97.8 / 99.5 / 94.7 98.4 / 99.4 / 94.4
Mean 79.1 / 91.3 / 59.1 85.5 / 95.9 / 75.6 87.2 / 96.8 / 81.4 86.8 / 96.0 / 75.2 90.5 / 98.2 / 85.1 92.3 / 98.4 / 85.5
4.3 Anomaly Detection on VisA
@@
Baseline. We further conduct experiments on the challenging VisA dataset. On this dataset, we train a unified model with images of all categories. For comprehensive comparisons, we have selected state-of-the-art models including DRAEM Zavrtanik et al. (2021), UniAD You et al. (2022), Simplenet Liu et al. (2023), ViTAD Zhang et al. (2023), and DiAD He et al. (2024).
Evaluation on VisA. As shown in Table 3, our method once gain achieves outstanding performance. Compared with the pioneering unified framework UniAD, we outperform it greatly by 6.8% and 3.5% in I-AUC and P-AUC, respectively. When compared with the pure ViT based model ViTAD, we outperform it in 10 out of 12 categories for detection and exceed it in 6 categories for localization. Specially, we improve upon it by 1.9%, 0.2% and 0.4% in I-AUC, P-AUC and PRO, respectively. DiAD is one diffusion-based framework for multi-class anomaly detection. We surpass it by 1.5% and 1.2% in I-AUC and P-AUC, respectively. The consistent improvements over the competitors validate the superiority of our method.
4.4 Anomaly Detection on MVTec-3D
Table 4: Anomaly detection results on MVTec-3D. denotes reproduced results using official code. Bold and underline indicate the best and second best, respectively.
Metric
DRÆM*
(ICCV 2021)
UniAD*
(NeurIPS 2022)
DiAD
(AAAI 2024)
ViTAD
(arXiv 2023)
Ours
I-AUC 65.2 78.9 84.6 78.7 84.9
P-AUC 93.2 96.5 96.4 98.0 98.0
PRO 55.0 88.1 87.8 91.3 92.9
Baseline. To further verify the superiority of the proposed method, we conduct experiments on the MVTec-3D dataset and compare it with state-of-the-art models including DRAEM Zavrtanik et al. (2021), UniAD You et al. (2022), and DiAD He et al. (2024).
@@ 4.4 Anomaly Detection on MVTec-3D
Evaluation on MVTec-3D. As illustrated in Table 4, our approach outperforms all others across every metric evaluated. Notably, when compared to UniAD, our model demonstrates a considerable advantage, with improvements of 6.0%, 1.5%, and 4.8% in the I-AUC, P-AUC, and PRO, respectively. Compared to the recent diffusion-based method DiAD, our approach has achieved significant improvements in the evaluation metrics, with enhancements of 0.3%, 1.6%, and 5.1% in the I-AUC, P-AUC, and PRO, respectively. ViTAD, a competitive method based on Vision Transformers, also performs well. However, our method still outperforms ViTAD in I-AUC and PRO, further highlighting its effectiveness. These results substantiate the superior performance of our method.
4.5 Ablation Study
@@
Backbone. To prove the generality of our method, we conduct comprehensive experiments using various backbones. As shown in Table 6, We employed two mainstream CNN-based backbones (Wide-Resnet-50He et al. (2015) and EfficientNet-b4Tan and Le (2020)) and two mainstream ViT-based backbones (CLIPRadford et al. (2021) and DINOCaron et al. (2021)) for our experiments. Compared to other mainstream anomaly detection frameworks, our method maintains competitive performance while using the same pre-trained backbones. Specifically, comparing to the ViTAD-CLIP, our method achieves a 23.4% improvement in I-AUC and a 14.2% improvement in P-AUC. Given the reconstruction-based strategy and plain Transformer architecture adopted by ViTAD, this significant enhancement can be attributed to the dynamic difficulty escalation and learnable updates provided by our attention mechanism, which effectively prevents training stagnation and ensures the robustness of anomaly detection. When facing SimpleNet with WideResNet50, our method outperforms with a 2.3% increase in I-AUC and a 4.5% increase in P-AUC. Furthermore, the consistent performance across both CNN and ViT-based architectures suggests that our method is not constrained by the specific characteristics of the backbone models. Instead, it effectively integrates with various feature extraction mechanisms, enhancing their anomaly detection capabilities. This adaptability is crucial for real-world applications where the choice of backbone may vary based on computational constraints or specific task requirements. In conclusion, the results fully demonstrate the generality of the proposed attention-guided perturbation method.
Multi-Layer Features. Table 8 systematically studies the impact of using multi-layer features fusion through four configurations: (1) Using only the deepest layer $f_{11}$achieves suboptimal 97.7% I-AUC and 97.8% P-AUC. This aligns with the fact that high-level features lose spatial details through complex neural networks, limiting precise anomaly location and local anomaly region perception. (2) Compared with only using the deepest feature, combining shallow layers features $f_{2},f_{5}$ boosts detection (98.1% I-AUC) yet degrades localization performance(97.0% P-AUC). We attribute this to low-level noise amplification. While early layers retain object boundaries and detailed information, their sensitivity to texture variations increases false positives in complex backgrounds. (3) Fusing the mid-level layers’ features($f_{8}$) and the deepest layers’ features ($f_{11}$) could improve the performance at image-level, compared with only utilizing the deepest layer features. This further proves that the shallower features have a better representation of the basic texture and contour of the sample to help the model achieve better image-level performance. 4) Overall, integrating both shallow and deep features yields optimal performance. This evidences our core hypothesis: Multi-layer fusion is non-additive but emergent. Just when features from both shallow and deep features interact cooperatively, the model simultaneously resolves different-grained anomaly analyses.
@@
Mean Teacher Decoder. Table 8 validates the effectiveness of the Mean Teacher decoder. Using the decoder of the reconstructed model to generate attention maps serves as the baseline. Introducing the Mean Teacher decoder results in an improvement of 0.3% for detection and 0.1% for localization, respectively. The experimental results demonstrate that the Mean Teacher decoder effectively mitigates the high variance of learnable attention weights during the initial training stages. Specifically, the Mean Teacher decoder supports a more stabilized attention-guided perturbation strategy, ensuring more stable weight updates in essential early training phases. This approach contrasts with the baseline decoder, where perturbations in initialization-sensitive regions (such as object boundaries or texture anomalies) induce unstable gradient updates that degrade detection consistency and localization precision.
Easy-to-Hard Noise Addition. As analyzed in Sec. 3.3, the hyperparameter $\alpha$ controls feature-level noise intensity in our noise addition curriculum.Table 7 explores the impact of noise intensity. Specially, we can observe that (1)Low noise intensity ($\alpha=0.5$) reduces the difficulty of decoder reconstruction. It is not conducive to modeling normal samples. This causes the model to fall behind in image-level evaluation (98.5% vs. 98.7%). (2) Excessive noise intensity($\alpha=1.5$)induces severe semantic distortion in feature representations, leading to performance degradation of 0.4% I-AUC and 0.1% P-AUC compared to the optimal configuration. (3) Appropriate noise intensity($\alpha=1$) enhances the model’s ability to capture normal distribution and enables smooth training, resulting in near-optimal performance. (4)The proposed easy-to-hard noise with suitable intensity helps the model capture the target distribution and achieve the best performance. Its advantage is that it dynamically adjusts the noise intensity, which facilitates the model to learn the de-noise task from easy to hard. The results support the advantages of this method.
Table 8: Ablation Studies on Multi-level Features and Mean Distillation. “M.T.D” indicates generating attention maps with mean distillation, while “w.o.” means without using mean distillation.
@@ 4.6 Efficiency Comparison with SoTAs
Table 9: Efficiency comparison of different methods
Method Learnable Parameters(M) FLOPs(G) FPS
DRÆM (ICCV 2021) 97.42 198.15 95.17
UniAD (NeurIPS 2022) 7.48 6.46 286.92
DeSTSeg (CVPR 2023) 32.37 30.67 386.76
SimpleNet (CVPR 2023) 3.94 18.34 63.66
Ours 4.42 5.63 326.65

### v3

@@ 4.2 Anomaly Detection on MVTec-AD
Bottle 99.2 / 99.1 / - 100.0 / 98.7 / 96.6 100.0 / 98.6 / - - / 99.2 / - 100.0 / 98.0 / - 100.0 / 98.9 / - 100.0 / 99.1 / 94.6
Cable 99.9 / 97.6 / - 95.0 / 97.4 / 91.0 99.5 / 98.4 / - - / 97.3 / - 94.8 / 96.8 / - 99.9 / 98.1 / - 100.0 / 98.7 / 93.9
Capsule 98.5 / 94.3 / - 96.3 / 98.7 / 95.8 98.1 / 98.8 / - - / 99.1 / - 97.7 / 98.9 / - 99.5 / 98.5 / - 97.9 / 98.3 / 90.4
Hazelnut 100.0 / 99.7 / - 99.9 / 98.9 / 95.5 100.0 / 98.7 / - - / 99.6 / - 100.0 / 97.9 / - 100.0 / 99.5 / - 100.0 / 99.0 / 94.4
Metal_nut 98.7 / 99.5 / - 100.0 / 97.3 / 92.3 100.0 / 98.4 / - - / 98.6 / - 100.0 / 98.8 / - 100.0 / 98.8 / - 100.0 / 97.0 / 91.5
Pill 98.9 / 97.6 / - 96.6 / 98.2 / 96.4 96.6 / 97.4 / - - / 98.7 / - 99.0 / 98.6 / - 98.1 / 97.9 / - 98.2 / 98.6 / 96.4
Screw 93.9 / 97.6 / - 97.0 / 99.6 / 98.2 98.1 / 99.4 / - - / 98.5 / - 98.2 / 99.3 / - 96.9 / 99.1 / - 97.2 / 99.3 / 96.7
Toothbrush 100.0 / 98.1 / - 99.5 / 99.1 / 94.5 100.0 / 98.7 / - - / 99.3 / - 99.7 / 98.5 / - 100.0 / 99.4 / - 97.2 / 98.9 / 90.3
Transistor 93.1 / 90.9 / - 96.7 / 92.5 / 78.0 100.0 / 96.3 / - - / 89.1 / - 100.0 / 97.6 / - 98.3 / 96.2 / - 99.5 / 97.3 / 91.3
Zipper 100.0 / 98.8 / - 98.5 / 98.2 / 95.4 99.4 / 98.8 / - - / 99.1 / - 99.9 / 98.9 / - 98.5 / 97.9 / - 99.3 / 97.4 / 93.0
Carpet 97.0 / 95.5 / - 98.9 / 98.9 / 97.0 98.7 / 99.0 / - - / 96.1 / - 99.7 / 98.2 / - 99.0 / 98.5 / - 100.0 / 99.2 / 95.5
Grid 99.9 / 99.7 / - 100.0 / 99.3 / 97.6 98.2 / 98.7 / - - / 99.1 / - 99.7 / 98.8 / - 100.0 / 99.6 / - 100.0 / 98.9 / 96.2
Leather 100.0 / 98.6 / - 100.0 / 99.4 / 99.1 100.0 / 99.3 / - - / 99.7 / - 100.0 / 99.2 / - 100.0 / 99.8 / - 100.0 / 99.5 / 97.7
Tile 99.6 / 99.2 / - 99.3 / 95.6 / 90.6 98.7 / 95.6 / - - / 98.0 / - 99.8 / 97.0 / - 100.0 / 98.7 / - 100.0 / 97.5 / 89.1
Wood 99.1 / 96.4 / - 99.2 / 95.3 / 90.9 99.2 / 95.0 / - - / 97.7 / - 100.0 / 94.5 / - 99.4 / 98.4 / - 98.8 / 96.7 / 88.5
Mean 98.0 / 97.3 / - 98.5 / 97.8 / 93.9 99.1 / 98.1 / 93.5 98.6 / 97.9 / - 99.6 / 98.1 / - 99.3 / 98.6 / - 99.2 / 98.3 / 93.1
For more details on each category, As shown in Table 3, our method achieves a leading performance even when trained on single-class normal data. It should be noted that our performance under multi-class setting is comparable to our multi-class results. However, existing methods like SimpleNet can attain near-saturation levels (e.g., 99.6% mean I-AUC) but with much inferior performance on the multi-class setting (see Table 1). It is also worth mentioning that our approach has achieved 100% I-AUC on 7 categories (Bottle, Hazelnut, Metalnut, Carpet, Grid, Leather, Tile), surpassing or matching state-of-the-art results. Based on the results, we can draw the following two conclusions:
1) On challenging texture categories like Carpet and Grid, we attain a perfect 100% I-AUC with 99.2% and 98.9% P-AUC, respectively. This outperforms SimpleNet (99.7%/98.2% for Carpet) and PatchCore (98.7%/98.7% for Grid), demonstrating that our attention-guided perturbation strategy effectively preserves structural consistency while amplifying subtle anomalies.
2) For objects requiring fine-grained defect localization (e.g., Screw: 97.2% I-AUC/99.3% P-AUC/96.7% PRO), our method significantly improves PRO scores by 1.4–5.2% over RD4AD and PatchCore. This highlights the advantage of dual-domain attention in distinguishing micro-defects from real anomalies.
Overall, the detailed experience results for each category are further evidence of the powerful ability of our attention-guided perturbation network.
Figure 4: Qualitative illustration on MVTec-AD dataset.
Figure 5: Qualitative illustration on Visa dataset.
4.3 Anomaly Detection on VisA and MVTec-3D
@@
Baseline. We further conduct experiments on the challenging VisA dataset under multi-class setting. For comprehensive comparisons, we have selected state-of-the-art models including DRAEM [3], UniAD [26], Simplenet [12], ViTAD [60] and DiAD [22]. On the MVTec-3D dataset, we compare our method with existing models, including DRAEM [3], UniAD [26], and DiAD [22].
Evaluation on VisA. As shown in Table 1, our method demonstrates competitive performance compared to existing approaches. Compared with the pioneering unified framework UniAD, we outperform it greatly by 6.8% and 3.5% in I-AUC and P-AUC, respectively. When compared with the pure ViT based model ViTAD, we outperform it in 10 out of 12 categories for detection and exceed it in 6 categories for localization. Specially, we improve upon it by 1.9%, 0.2% and 0.4% in I-AUC, P-AUC, and PRO, respectively. DiAD is one diffusion-based framework for multi-class anomaly detection. We surpass it by 1.5% and 1.2% in I-AUC and P-AUC, respectively. While our method exhibits marginally lower performance compared to the state-of-the-art MoEAD on VisA (e.g., 98.4% vs. 98.7% P-AUC), the experimental results across multiple datasets(e.g., MVTec AD 98.0% vs. 97.0% P-AUC) demonstrate its consistent effectiveness. Figure 5 illustrates our method’s localization performance on the Visa dataset. Rows 3 and 8 demonstrate precise detection of microscale anomalies, enabled by our multiscale feature integration that preserves fine-grained spatial details. The heatmap in Row 2, Column 3 highlights robust localization of significant structural anomalies, with activation maps tightly aligning to anomaly region boundaries.
@@
Evaluation on MVTec-3D. As illustrated in Table 1, our approach outperforms all others across every metric evaluated. When compared to UniAD, our model demonstrates a considerable advantage, with improvements of 6.0%, 1.5%, and 4.8% in the I-AUC, P-AUC, and PRO, respectively. Compared to the diffusion-based method DiAD, our approach has achieved significant improvements in the evaluation metrics, with enhancements of 0.3%, 1.6%, and 5.1% in the I-AUC, P-AUC, and PRO, respectively. ViTAD, a competitive method based on Vision Transformers, also performs well. However, our method still outperforms ViTAD in I-AUC and PRO, further highlighting its effectiveness. These results substantiate the superior performance of our method.
@@ 4.4 Ablation Study
Backbone. To prove the generality of our method, we conduct comprehensive experiments using various backbones. As shown in 5, We employed two mainstream CNN-based backbones (Wide-Resnet-50[63] and EfficientNet-b4[64]) and two mainstream ViT-based backbones (CLIP[50] and DINO[53]) for our experiments. Compared to other mainstream anomaly detection frameworks, our method maintains competitive performance while using the same pre-trained backbones. Specifically, comparing to the ViTAD-CLIP, our method achieves a 23.4% improvement in I-AUC and a 14.2% improvement in P-AUC. When facing SimpleNet with WideResNet50, our method outperform with a 2.3% increase in I-AUC and a 4.5% increase in P-AUC. The results fully demonstrate the generality of the proposed attention-guided perturbation method.
Multi-Layer Features.  7 studies the impact of using multi-layer features. It can be observed that when utilizing deep features, the model demonstrates robust detection capacities, achieving 97.8% I-AUC. In contrast, employing shallow features exhibits desirable localization abilities. Overall, integrating both shallow and deep features yields optimal performance.
@@
Mean Teacher Decoder. Table 7 validates the effectiveness of the Mean Teacher decoder. Using the decoder of the reconstructed model to generate attention maps serves as the baseline. Introducing the Mean Teacher decoder results in an improvement of 0.3% for detection and 0.1% for localization. The experimental results demonstrate that the Mean Teacher decoder effectively mitigates the high variance of learnable attention weights during the initial training stages.
Easy-to-Hard Noise Addition. Recall that in Sec. 3.3, we use a hyperparameter $\alpha$ to control the noise intensity at the feature level. 6 explores the impact of noise intensity. The experiments indicate that Easy-to-hard noise with suitable intensity helps the model to model the target distribution and achieve the best performance. Compared with fixed noise of different intensities, the I-AUC and P-AUC are slightly improved(0.1%-0.4%).
Table 7: Ablation Studies on Multi-level Features and Mean Distillation. “M.T.D” indicates generating attention maps with mean distillation, while “w.o.” means without mean distillation.
@@
Table 8: Efficiency comparison of different methods
Method Learnable Parameters FLOPs FPS
DRÆM (ICCV 2021) 97.42M 198.15G 95.17
UniAD (NeurIPS 2022) 7.48M 6.46G 286.92
DeSTSeg (CVPR 2023) 32.37M 30.67G 386.76
SimpleNet (CVPR 2023) 3.94M 18.34G 63.66
DiAD (AAAI 2024) 1.33B ¿2.2T -
Ours 4.42M 5.63G 326.65